July 18, 2005
As I mentioned in yesterday's entry, the incubation experiments have kept Caroline and Sean very busy. Julia and Maureen have helped them filter their experiments as I am still filtering samples from the CTD (casts 98-113). Today we completed two lines: CB (Copalis Beach) and KB (Kalaloch Beach). Fortunately, Brian and James were able to relieve me at the end of the day.
Besides the light not working on the EH3 mooring, there have been other problems to contend with. Earlier today, the Lachat nutrient analyzer that is used by Julian Herndon started to act up. So Julian worked on the analyzer all day trying to figure what was wrong. He and Eric opened up the analyzer to inspect the wiring and instrumentation. Their initial assessment revealed that it had to be something wrong with the chemistry taking place within the instrument. Julian spent the rest of the night looking through manuals and emailing the Lachat Company in for troubleshooting tips. For the time being, I will now have to freeze all of the nutrient samples collected from the CTD. The worst case scenario is that Julian will have to run the samples at the SFSU lab after the cruise. If it is fixed within the next few days, I can continue to collect nutrient samples and place them in the refrigerator rather than freeze them.
That was not the only problem of the day. One of the recirculating water baths used in the rad-van to maintain the temperature of P vs. E incubators malfunctioned. P vs. E isotopic experiments are performed to determine the photosynthetic efficiency of phytoplankton. The samples are exposed to different amounts of light to determine their efficiency of carbon-14 uptake. This will make it harder for Julia, Maureen, Liza and Dr. Cochlan to conduct their experiments. The ship's engineers will try to fix the water bath tomorrow.